OpenStudy (anonymous):

detrmine3/4 of 20,64,36

thomaster (thomaster):

\(\Huge◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►\\\huge\cal\color{red}{~~~~~~~Welcome\ to\ OpenStudy!!~~ツ}\\\Huge◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►◄►\) Do you mean what the \(\Large\frac{3}{4}\) part of 20, 64 and 36 is?

OpenStudy (anonymous):

yes

thomaster (thomaster):

Oke \(\huge \frac{3}{4}\LARGE*20\) is the same as \(\huge \frac{3}{4}*\frac{20}{1}\) To solve that you do \(\huge \frac{3*20}{4*1}~\to~\frac{60}{4}\) Do you know the rest?

thomaster (thomaster):

@savitarawat are you still there?